OpenStudy (anonymous):

Find the odds. The kings are separated from a deck of standard playing cards and shuffled. One king is randomly selected. What are the odds in favor of drawing a black card? A. 4 to 1 B. 1 to 4 C. 1 to 2 D. 1 to 1

OpenStudy (anonymous):

I can help...hold on one second

OpenStudy (anonymous):

P(black)=26/52 or 1/2, P(King) is 42/52 or 1/13 so P(black and king) = 1/2*1/13 = 1/26